# Answering the Unanswerable Is to Err Knowingly: Analyzing and Mitigating Abstention Failures in Large Reasoning Models

### v1

@@ 1 Introduction
Most current research on LRMs focuses on improving the reasoning process for solving complex problems (Xu et al. 2025), such as reducing redundant reasoning steps to alleviate overthinking (Dang, Huang, and Chen 2025; Yang et al. 2025b) and refining the reasoning process to further enhance reasoning performance (Chen et al. 2025; Liao et al. 2025). However, limited work has considered that not all complex questions are answerable (Kirichenko et al. 2025; Song, Shi, and Zhao 2025). When confronted with noisy and diverse user queries, there will always be cases where LRMs are unable to produce reliable responses, such as the math problems missing sufficient conditions.
As illustrated in Figure 1, when a question is unanswerable due to missing necessary conditions, we prefer LRMs to reason about why it cannot be answered and have the ability to abstain by responding with “I don’t know” (i.e., abstention success), rather than generating a reasoning process with hallucination and arriving at an incorrect answer (i.e., abstention failure). Since model reliability is foundational to user trust, LRMs need to possess both strong reasoning abilities and the capacity to abstain from answering unanswerable questions (Kirichenko et al. 2025; Song, Shi, and Zhao 2025). We first validate the phenomenon that LRMs often struggle to abstain. As shown in Figure 2, we evaluate several LRMs on SUM (Song, Shi, and Zhao 2025), which contains mathematical unanswerable questions. Our results show that most LRMs fail to abstain on more than half of the examples. To address the above issue, we systematically analyze, investigate, and propose solutions to improve the abstention behavior for LRMs.
@@
Finally, we conduct extensive experiments with two datasets using LRMs from various model families and scales. Our method enhances LRMs’ ability to abstain from answering unanswerable questions without degrading their reasoning on answerable ones. Furthermore, our experiments reveal that different types of abstention failures benefit from different intervention strategies. Our method achieves significant improvements across all failure types. We will open-source our code to facilitate future research.
@@ 2 Related Work
As the demand for more reliable language models grows, the ability of LLMs to abstain from answering unanswerable questions has become an important evaluation criterion (Yin et al. 2023; Amayuelas et al. 2024; Madhusudhan et al. 2025; Sun et al. 2024; Tomani et al. 2024). LRMs have attracted attention for their strong performance on reasoning tasks (Yang et al. 2025b; Fu et al. 2025). However, their behavior on unanswerable questions has been less studied. Prior work (Kirichenko et al. 2025; Song, Shi, and Zhao 2025) finds that LRMs show weaker abstention ability. In our work, we further provide a more detailed analysis of the different types of outputs produced by LRMs when faced with unanswerable questions. We show that LRMs often have an internal understanding of a problem’s solvability but fail to express it through explicit abstention. Based on this insight, we propose a method to improve their abstention behavior.
@@
While LRMs benefit from richer chain-of-thought (CoT) reasoning and achieve strong performance on complex tasks, recent efforts have explored inference-time interventions to further enhance their reasoning accuracy and efficiency (Chen et al. 2025; Dang, Huang, and Chen 2025). For example, (Fu et al. 2025) and (Li et al. 2024) monitor intermediate consistency between reasoning steps to decide when to output the answer. (Yang et al. 2025b) estimates confidence in intermediate steps to decide whether the model has reached a sufficiently certain conclusion. (Liao et al. 2025) uses process-level reward models to guide the model toward better reasoning trajectories. However, the above methods are designed for answerable questions. In our work, we evaluate their effectiveness on unanswerable questions and propose a new inference-time intervention to improve the abstention capability for LRMs.
@@ 3 Analysis of Abstention Failure
In this section, we begin by analyzing how LRMs respond to unanswerable math problems. We then investigate LRMs’ awareness of unanswerable questions by examining whether they possess the ability to recognize the unanswerability of such questions, from both internal and external perspectives.
@@
Following previous work (Song, Shi, and Zhao 2025; Ouyang 2025), we focus on unanswerable math problems, i.e., ill-posed problems that are challenging for LRMs and offer objectively defined criteria for unanswerability. Specifically, we use the Synthetic Unanswerable Math (SUM) (Song, Shi, and Zhao 2025) dataset, which includes diverse problems from AIME (1984–2023), AMC (pre-2023), Omni-MATH (Gao et al. 2025), and Still (Team 2025). The unanswerable problems in SUM are generated based on five criteria: (1) key information deletion, (2) ambiguous key information, (3) unrealistic conditions, (4) unrelated objects, and (5) question deletion. We randomly sample 1,000 problems for our analysis. For each unanswerable case, we use GPT-4o to generate a brief explanation as the ground-truth rationale.
@@ 3.1 Reactions to Unanswerable Questions
In this section, we evaluate the reactions of LRMs to unanswerable questions using the prompts in Figure 3 with a maximum token budget of 10,000 per response. We identify three distinct response types, as shown in Figures 4:
@@ 3.2 Awareness of Unanswerable Questions
Figure 7: The proportion of questions that can provide a correct explanation in stopping points during the reasoning in the types of “hallucinated answer” and “cognitive fixation”.
@@
These findings suggest that even when an LRM fails to abstain, it may still recognize unanswerability during reasoning. Its ability to assess answerability exists but is underused in final decisions.
@@
We conduct a probing-based analysis on the latent representations during reasoning. The linear representation hypothesis posits that high-level concepts, such as language, gender, and truthfulness, are linearly embedded in the latent space of LLMs  (Park, Choe, and Veitch 2024). We hypothesize that question answerability may also be linearly represented during the reasoning process. Inspired by prior work on representation probing (Li et al. 2023; Orgad et al. 2025; Marks and Tegmark 2023; Slobodkin et al. 2023), we train lightweight linear classifiers (probes) on hidden activations from the LRMs’ reasoning trajectory. The goal is to assess whether answerable and unanswerable questions can be distinguished from latent representations during the reasoning process, thereby revealing the presence of answerability-related signals in internal state.
@@ 4 Mitigation of Abstention Failure
In this section, we seek to mitigate abstention failures of LRMs when faced with unanswerable questions.
4.1 Motivation
As previously discussed, although LRMs show signs of recognizing unanswerable questions during reasoning, they often fail to abstain in their final answers. We aim to identify the causes of this misalignment and explore ways to fix it.
First, we examine the LRMs’ confidence (Kuhn, Gal, and Farquhar 2023) in abstaining at the stopping point across three output types. We sample an equal number of questions from each output type and compute the average confidence in generating “I don’t know” at the stopping points. Following (Yang et al. 2025b), the confidence score $C$ is computed as
$\begin{aligned} p(a_{t})=P(a_{t}\,|\,H,I,a_{<t}),C=\left(\sum_{i=1}^{n}\max_{a_{t}\in V}p(a_{t})\right)^{\frac{1}{n}},\end{aligned}$ (2)
where $p(a_{t})$ denotes the model’s predicted probability of answer token $a_{t}$, $H$ includes the input prompt and the generated thoughts, $I$ is the prompt used to elicit the answer, $n$ is the number of decoding steps, and $V$ is the vocabulary. As shown in Figure 9, the types of hallucinated answer and cognitive fixation both exhibit lower confidence in abstention compared to the type of correct abstention.
@@ 4.1 Motivation
Next, we analyze the average frequency of “I don’t know” outputs by the LRMs across all stopping points in the reasoning trajectory. As shown in Figure 10, the hallucinated answer and cognitive fixation types consistently show lower abstention frequencies than correct abstention.
@@ 4.2 Method
The first step of our method aims to identify when the model internally recognizes that a question may be unanswerable. To do so, we track the token-level hidden states generated during inference and segment the reasoning process into semantically coherent units (e.g., clauses or transitions marked by discourse cues such as “wait”). At the end of each segment, we apply a lightweight linear probe, which is trained from our analysis of latent signals of question answerability, to estimate the probability that the current question is unanswerable. If the predicted probability of unanswerability exceeds a threshold, the model transitions to the next stage: inference-time intervention.
@@ 5.1 Setup
We use SUM  (Song, Shi, and Zhao 2025) and UMWP  (Sun et al. 2024). For SUM, we used its test set, which contains 284 unanswerable and 284 answerable questions manually verified. UMWP is derived from four widely used math word problem datasets: SVAMP (Patel, Bhattamishra, and Goyal 2021), MultiArith (Koncel-Kedziorski et al. 2016), GSM8K (Cobbe et al. 2021), and ASDiv (Miao, Liang, and Su 2020). The full dataset consists of 5,200 instances. We sample 600 questions (300 unanswerable and 300 answerable) to form the test set. For the unanswerable questions in UMWP, we also use GPT-4o to generate the reasoning behind their unanswerability as ground truth.
@@
Following previous works (Song, Shi, and Zhao 2025; Kirichenko et al. 2025), we evaluate model performance using the following metrics. (1) Abstention Rate: The proportion of unanswerable questions for which the model correctly abstains from answering by outputting “I don’t know” in accordance with the instruction. (2) Reason Accuracy: The percentage of unanswerable questions for which the model provides the correct rationale for unanswerability. (3) Token Usage: The number of tokens generated by the model during the reasoning process. (4) Answer Accuracy: The proportion of answerable questions for which the model produces the correct final answer.
@@
Our method is an inference-time intervention for LRMs without training. There are no directly comparable baselines. We select two baselines based on output consistency (Dynasor-CoT) and confidence (DEER), assuming that correct abstention is the correct answer for unanswerable questions. Dynasor-CoT (Fu et al. 2025) prompts intermediate answers and halts when the same answer appears three times consecutively. DEER (Yang et al. 2025b) monitors sentence-level confidence and exits early once a threshold is met. Besides, the Vanilla method denotes unaltered LRM outputs. For LRMs, we choose five models: R1-Distill-Llama-8B, R1-Distill-Qwen-7B, R1-Distill-Qwen-14B, Qwen3-8B, and Qwen3-14B. Implementation details and full results for all datasets and LRMs are in the appendix.
@@ 5.2 Main Results
Table 1 shows the main results. We have the following observations: (1) Our method achieves the highest Abstention Rate and Reason Accuracy on unanswerable questions, demonstrating its strong ability to guide LRMs in recognizing and abstaining from unanswerable inputs. (2) Our method maintains comparable Answer Accuracy on answerable questions. In most settings, accuracy remains close to the vanilla model and sometimes improves slightly, indicating minimal impact on solving answerable tasks. (3) Our method can reduce token usage on unanswerable questions, with an average reduction of 30–50% across different LRMs compared to the vanilla model. It also slightly decreases token consumption on answerable questions, indicating improved reasoning efficiency. (4) We observe a positive correlation between Abstention Rate and Reason Accuracy: as the LRM becomes better at abstaining from unanswerable questions, it also produces more accurate explanations. This suggests that the intervention not only changes the final output but also improves intermediate reasoning quality. (5) Qwen3 models generally outperform distillation-based models in abstention-related metrics. Larger LRMs tend to exhibit stronger abstention capabilities. This trend indicates that both model scale and architecture contribute to reliable unanswerability detection.
@@ 5.3 Further Discussions
We analyze the changes in the proportions of different response types across methods and datasets, as shown in Figure 12. Our method consistently reduces the hallucinated answers and cognitive fixation outputs. This reduction directly contributes to a substantial increase in the rate of correct abstentions. While both Dynasor-CoT and DEER employ early-exit strategies to mitigate cognitive fixation, we observe that such baselines often lead to a higher proportion of hallucinated answers. Early exits without appropriate guidance may cause the model to make up assumptions or imagined scenarios for giving a definite answer, rather than acknowledging uncertainty. This highlights the importance of combining monitoring with guided intervention to steer LRMs toward proper behavior.
@@
We assess how our intervention influences the LRM’s confidence in abstention and its actual abstention behavior, in order to evaluate whether our method helps bridge the gap between cognitive awareness and abstention behavior. Specifically, at the intervention point identified by our method, we prompt the LRM to generate intermediate outputs both before and after the intervention. We then measure two key indicators: the confidence when producing “I don’t know” responses, and the proportion of questions for which the model outputs “I don’t know”. The results are shown in Table 2. Our method consistently enhances the confidence in generating abstention responses. In addition, the abstention rate also shows corresponding improvements.
@@
We further analyze the cognitive monitoring component of our method and compare our default monitoring strategy based on latent representations with alternative strategies relying on behavioral signals. The behavioral signal approach monitors the LRMs’ intermediate outputs at the end of the paragraph generation phase (e.g., when it reaches a “wait” token), and uses these outputs to determine whether to trigger an intervention. We investigate three variants: The Direct Behavior strategy checks whether the model’s intermediate output is “I don’t know” and triggers an intervention immediately if so. The Consistency strategy triggers intervention only if the model produces “I don’t know” in three consecutive intermediate outputs (inspired by Dynasor-CoT). The Confidence Score strategy triggers intervention when the model outputs “I don’t know” with a confidence score exceeding a predefined threshold (inspired by DEER). As shown in Table 3, all monitoring strategies contribute to the improvement in abstention behavior, showing that cognitive monitoring is generally effective. Among them, the strategy based on latent representation signals achieves the best and most consistent performance across different models and datasets. The Direct Behavior method is simple and works well, but it can be too aggressive and may hurt performance on answerable questions.
@@
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
@@
We evaluate two aspects of inference-time intervention: the instructional guidance prompt and the early exit strategy. By removing one component at a time, we analyze how each affects abstention behavior and answer quality. The results are shown in Table 4. For correct abstention, the impact of instructional guidance is greater than that of early exit. The early exit strategy helps reduce the number of cognitive fixation cases. However, without instructional guidance, the proportion of hallucinated answers increases. This again shows that without proper guidance, the model tends to make up conditions and generate unsupported answers. Instructional guidance also has a slight impact on the performance of answerable questions.
@@
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$

### v3

@@ 1 Introduction
Most current research on LRMs aims to improve the reasoning process for solving complex problems (Xu et al. 2025), such as reducing redundant reasoning steps to alleviate overthinking (Yang et al. 2025b) and refining the reasoning process to further enhance reasoning performance (Chen et al. 2025; Liao et al. 2025). Limited work has considered that not all complex questions are answerable (Kirichenko et al. 2025; Song et al. 2025). When confronted with noisy and diverse user queries, there will always be cases where LRMs are unable to produce reliable responses, such as the math problems missing sufficient conditions.
In Figure 1, when a question is unanswerable due to missing necessary conditions, we prefer LRMs to reason about why it cannot be answered and have the ability to abstain by responding with “I don’t know” (i.e., abstention success), rather than generating a reasoning process with hallucination and arriving at an incorrect answer (i.e., abstention failure). Since model reliability is foundational to user trust, LRMs need to possess both strong reasoning abilities and the capacity to abstain from answering unanswerable questions (Kirichenko et al. 2025; Song et al. 2025). We first validate the phenomenon that LRMs often struggle to abstain. In Figure 2, we evaluate several LRMs on SUM (Song et al. 2025), which contains mathematical unanswerable questions. Our results show that most LRMs fail to abstain on more than half of the examples. To address the above issue, we systematically analyze, investigate, and propose solutions to improve the abstention behavior for LRMs.
@@
Finally, we conduct extensive experiments with two datasets using LRMs from various model families and scales. Our method enhances LRMs’ ability to abstain from answering unanswerable questions without degrading their reasoning on answerable ones. Furthermore, our experiments reveal that different types of abstention failures benefit from different intervention strategies. Our method achieves significant improvements across all failure types. We release our code at https://github.com/nju-websoft/AbstentionReasoning.
@@ 2 Related Work
As the demand for more reliable language models grows, the ability of LLMs to abstain from answering unanswerable questions has become an important evaluation criterion (Yin et al. 2023; Amayuelas et al. 2024; Madhusudhan et al. 2025; Sun et al. 2024; Tomani et al. 2024). LRMs have attracted attention for their strong performance on reasoning tasks (Yang et al. 2025b; Fu et al. 2025). However, their behavior on unanswerable questions has been less studied. Prior work (Kirichenko et al. 2025; Song et al. 2025) finds that LRMs show weaker abstention ability. We further provide a more detailed analysis of the different types of outputs produced by LRMs when faced with unanswerable questions. We show that LRMs have an internal understanding of a problem’s solvability but fail to express it through explicit abstention, and propose a method to improve their abstention behavior.
@@
While LRMs benefit from richer chain-of-thought (CoT) reasoning and achieve strong performance on complex tasks, recent efforts have have begun to analyze (Dang et al. 2025; Zhu et al. 2025; Liu et al. 2025) and implement (Chen et al. 2025) inference-time interventions to further enhance their reasoning accuracy and efficiency. For example, (Fu et al. 2025) and (Li et al. 2024) monitor intermediate consistency between reasoning steps to decide when to output the answer. (Yang et al. 2025b) estimates confidence in intermediate steps to decide whether the model has reached a sufficiently certain conclusion. (Liao et al. 2025) uses process-level reward models to guide the model toward better reasoning trajectories. However, the above methods are designed for answerable questions. In our work, we evaluate their effectiveness on unanswerable questions and propose a new inference-time intervention to improve the abstention capability for LRMs.
@@ 3 Analysis of Abstention Failure
We first analyze how LRMs respond to unanswerable math problems. Then, we investigate LRMs’ awareness of unanswerable questions by examining whether they possess the ability to recognize the unanswerability of such questions, from both internal and external perspectives.
@@
Following previous work (Song et al. 2025; Ouyang 2025), we focus on unanswerable math problems, i.e., ill-posed problems that are challenging for LRMs and offer objectively defined criteria for unanswerability. Specifically, we use the Synthetic Unanswerable Math (SUM) (Song et al. 2025) dataset, which includes diverse problems from AIME (1984–2023), AMC (pre-2023), Omni-MATH (Gao et al. 2025), and Still (Team 2025). The unanswerable problems in SUM are generated based on five criteria: (1) key information deletion, (2) ambiguous key information, (3) unrealistic conditions, (4) unrelated objects, and (5) question deletion. We randomly sample 1,000 problems for our analysis. For each unanswerable case, we use GPT-4o to generate a brief explanation as the ground-truth rationale.
@@ 3.1 Reactions to Unanswerable Questions
We evaluate the reactions of LRMs to unanswerable questions using the prompts in Figure 3 with a maximum token budget of 10,000 per response. We identify three distinct response types, as shown in Figures 4:
@@ 3.2 Awareness of Unanswerable Questions
Figure 7: The proportion of questions that can provide a correct explanation in stopping points during the reasoning.
@@
Overall, even when an LRM fails to abstain, it may still recognize unanswerability during reasoning. Its ability to assess answerability exists but is underused in final decisions.
@@
We conduct a probing-based analysis on the latent representations during reasoning. The linear representation hypothesis posits that high-level concepts, such as language, gender, and truthfulness, are linearly embedded in the latent space of LLMs  (Park et al. 2024). We hypothesize that question answerability may also be linearly represented during the reasoning process. Inspired by prior work on representation probing (Li et al. 2023; Orgad et al. 2025; Marks and Tegmark 2023; Slobodkin et al. 2023), we train lightweight linear classifiers (probes) on hidden activations from the LRMs’ reasoning trajectory. The goal is to assess whether answerable and unanswerable questions can be distinguished from latent representations during the reasoning process, thereby revealing the presence of answerability-related signals in internal state.
@@ 4 Mitigation of Abstention Failure
We seek to mitigate abstention failures of LRMs when faced with unanswerable questions. Although LRMs show signs of recognizing unanswerable questions during reasoning, they often fail to abstain in their final answers. We aim to identify the causes of this misalignment and explore ways to fix it.
First, we examine the LRMs’ confidence (Kuhn et al. 2023) in abstaining at the stopping point across three output types. We sample an equal number of questions from each output type and compute the average confidence in generating “I don’t know” at the stopping points. Following (Yang et al. 2025b), the confidence score $C$ is computed as
$\begin{aligned} p(a_{t})=P(a_{t}\,|\,H,I,a_{<t}),C=\left(\prod_{i=1}^{n}\max_{a_{t}\in V}p(a_{t})\right)^{\frac{1}{n}},\end{aligned}$ (2)
where $p(a_{t})$ is the model’s predicted probability of answer token $a_{t}$, $H$ includes the input prompt and the generated thoughts, $I$ is the prompt to elicit the answer, $n$ is the number of decoding steps, and $V$ is the vocabulary. In Figure 9, hallucinated answer and cognitive fixation both exhibit lower confidence in abstention compared to correct abstention.
@@
Next, we analyze the average frequency of “I don’t know” outputs by the LRMs across all stopping points in the reasoning trajectory. In Figure 10, the hallucinated answer and cognitive fixation types consistently show lower abstention frequencies than correct abstention.
@@
The first step of our method aims to identify when the model internally recognizes that a question may be unanswerable. We track the token-level hidden states generated during inference and segment the reasoning process into semantically coherent units (e.g., clauses or transitions marked by discourse cues such as “wait”). At the end of each segment, we apply a lightweight linear probe, which is trained from our analysis of latent signals of question answerability, to estimate the probability that the current question is unanswerable. If the predicted probability exceeds a threshold, the model takes the intervention strategy.
@@ 5.1 Setup
We use SUM  (Song et al. 2025) and UMWP  (Sun et al. 2024). For SUM, we used its test set, which contains 284 unanswerable and 284 answerable questions manually verified. UMWP is derived from four widely used math word problem datasets: SVAMP (Patel et al. 2021), MultiArith (Koncel-Kedziorski et al. 2016), GSM8K (Cobbe et al. 2021), and ASDiv (Miao et al. 2020). The full dataset consists of 5,200 instances. We sample 600 questions (300 unanswerable and 300 answerable) to form the test set. For the unanswerable questions in UMWP, we also use GPT-4o to generate the reasoning behind their unanswerability as ground truth.
@@
Following (Song et al. 2025; Kirichenko et al. 2025), we evaluate model performance with (1) Abstention Rate: The proportion of unanswerable questions for which the model correctly abstains from answering by outputting “I don’t know” in accordance with the instruction. (2) Reason Accuracy: The percentage of unanswerable questions for which the model provides the correct rationale for unanswerability. (3) Token Usage: The number of tokens generated by the model during the reasoning process. (4) Answer Accuracy: The proportion of answerable questions for which the model produces the correct final answer.
@@
Our method is an inference-time intervention for LRMs without training. There are no directly comparable baselines. We select two baselines based on output consistency (Dynasor-CoT) and confidence (DEER), assuming that correct abstention is the correct answer for unanswerable questions. Dynasor-CoT (Fu et al. 2025) prompts intermediate answers and halts when the same answer appears three times consecutively. DEER (Yang et al. 2025b) monitors sentence-level confidence and exits early once a threshold is met. Besides, the Vanilla method denotes unaltered LRM outputs. For LRMs, we choose R1-Distill-Llama-8B, R1-Distill-Qwen-7B, R1-Distill-Qwen-14B, Qwen3-8B, and Qwen3-14B. More details are in the appendix.
@@ 5.2 Main Results
Table 1 shows the main results. We have the following observations: (1) Our method achieves the highest Abstention Rate and Reason Accuracy on unanswerable questions, demonstrating its strong ability to guide LRMs in recognizing and abstaining from unanswerable inputs. (2) Our method maintains comparable Answer Accuracy on answerable questions. In most settings, accuracy remains close to the vanilla model and sometimes improves slightly, indicating minimal impact on solving answerable tasks. (3) Our method can reduce token usage on unanswerable questions, with an average reduction of 30–50% across different LRMs compared to the vanilla model. It also slightly decreases token consumption on answerable questions, indicating improved reasoning efficiency. (4) We observe a positive correlation between Abstention Rate and Reason Accuracy: as the LRM becomes better at abstaining from unanswerable questions, it also produces more accurate explanations. This suggests that the intervention not only changes the final output but also improves intermediate reasoning quality. (5) Qwen3 models generally outperform other models in abstention-related metrics. Larger LRMs tend to exhibit stronger abstention capabilities. This trend indicates that both model scale and architecture contribute to reliable unanswerability detection.
@@ 5.3 Further Discussions
We analyze the changes in the proportions of different response types across methods and datasets. Figure 12 shows that our method consistently reduces the hallucinated answers and cognitive fixation outputs, which contribute to a substantial increase in the rate of correct abstentions. While Dynasor-CoT and DEER employ early exits to mitigate cognitive fixation, they lead to a higher proportion of hallucinated answers. Early exits without appropriate guidance may cause the model to make up assumptions or imagined scenarios for giving a definite answer, rather than acknowledging uncertainty. This highlights the importance of combining monitoring with guided intervention to steer LRMs toward proper behavior.
@@
We assess how our intervention influences the LRM’s confidence in abstention and its actual abstention behavior, to evaluate whether our method helps bridge the gap between cognitive awareness and abstention behavior. At the intervention point identified by our method, we prompt the LRM to generate intermediate outputs before and after the intervention. We measure two key indicators: the confidence when producing “I don’t know” responses, and the proportion of questions for which the model outputs “I don’t know”. Table 2 shows that our method consistently enhances the confidence in generating abstention responses. The abstention rate also shows corresponding improvements.
@@
We analyze the cognitive monitoring component and compare our default monitoring strategy based on latent representations with alternative strategies relying on behavioral signals. The behavioral signal approach monitors the LRMs’ intermediate outputs at the end of the paragraph generation phase (e.g., when it reaches a “wait” token), and uses these outputs to determine whether to trigger an intervention. We investigate three variants: The Direct Behavior strategy checks whether the model’s intermediate output is “I don’t know” and triggers an intervention immediately if so. The Consistency strategy triggers intervention only if the model produces “I don’t know” in three consecutive intermediate outputs (inspired by Dynasor-CoT). The Confidence Score strategy triggers intervention when the model outputs “I don’t know” with a confidence score exceeding a predefined threshold (inspired by DEER). In Table 3, all monitoring strategies contribute to the improvement in abstention behavior, showing that cognitive monitoring is generally effective. The strategy based on latent representation signals achieves the best and most consistent performance across models and datasets. The Direct Behavior method is simple and works well, but it may be too aggressive and hurt performance on answerable questions.
@@
Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$
@@
We evaluate two aspects of inference-time intervention: the instructional guidance prompt and the early exit strategy. We analyze how each aspect affects abstention behavior and answer quality. The results are shown in Table 4. For correct abstention, the impact of instructional guidance is greater than that of early exit. The early exit strategy helps reduce the number of cognitive fixation cases. Without instructional guidance, the proportion of hallucinated answers increases. This again shows that without proper guidance, the model tends to make up conditions and generate unsupported answers. Instructional guidance also has a slight impact on the performance of answerable questions.
@@
Correct abstention $\uparrow$ Hallucinated answer $\downarrow$ Cognitive fixation $\downarrow$ Acc $\uparrow$